OpenStudy (anonymous):

not understanding how to do this. can someone help? Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). A. y = -4x + 1 B. y = -1/2 x + 9 C. y = 4x + 1 D. y = -2x + 9

OpenStudy (anonymous):

1) Find slope. \[Slope = \frac{ y2 - y1 }{ x2 - x1 }\]\[Slope = \frac{ -1 - 1 }{ 5 - 4 } = \frac{ -2 }{ 1 } = -2\] So you have a slope of -2. You also have a given point. 2) Put into Point-Slope form. \[y - y1 = m(x - x1) \] Lets use point (4, 1) and the slope of -2. \[y - 1 = -2(x - 4)\]3) Change into Slope-Intercept form. \[y - 1 = -2(x - 4)\]Distribute.\[y - 1 = -2x + 8\]Add 1 to both sides.\[y = -2x + 9\] Slope-intercept form is: \[y = -2x + 9\]

OpenStudy (anonymous):

oooh thank you so much for showing me how to do this. i think im getting to know how to do it better now. thank you again!

OpenStudy (anonymous):

You're welcome! I'm glad this helped you!(: Also:\[ \Huge{\color{turquoise}{\mathbb{Welcome \quad To \quad OpenStudy!!}}} \]

OpenStudy (anonymous):

yes it truly did help. thank you